# Implementation and characterization of the dice lattice in the electron quantum simulator

### v2

@@ III Conclusions and Outlook
In this study, we explored the realization of a dice lattice on a Cu(111) surface decorated with CO molecules. Our experimental proposal demonstrates that the key spectral features of the dice lattice are maintained in some of the proposed implementations. We have verified this by fitting the electronic spectral properties using a next-nearest-neighbor tight-binding model. Some of the proposed implementations reveal a touching point and a quasi-flat band.
@@
As a final perspective for our results, we envision two important developments: as for the case of the honeycomb lattice, an opportune modulation of the lattice periodicity can result in a homogeneous shift of the Fermi energy [18]. This property can use used to create an effective $pn$-juntion where the super-Klein tunneling [42] could be observed. Secondly, to enhance the possibility of observing the effects of electron correlation in the flat band, we plan to investigate in future studies the implementation of the dice lattice in semiconducting electron simulator platforms based on InAs and InSb decorated with In or Cs adatoms [54, 55].

### v3

@@ III Conclusions and Outlook
In this study, we explored the realization of a dice lattice on a Cu(111) surface decorated with CO molecules. Our experimental proposal demonstrates that the key spectral features of the dice lattice are maintained in some of the proposed implementations. We have verified this by fitting the electronic spectral properties using a next-nearest-neighbor tight-binding model. Some of the proposed implementations reveal a touching point and a quasi-flat band. We note on passing that one largest experiments performed with the electron quantum simulator for the realization of the Pensore tiling [21] included around 460 CO molecules. Extrapolating on this number would mean we could achieve around 20 unit cells for the dice lattice in the CII configuration. This number of unit cells is sufficient to already observe bulk properties in finite-size systems.
@@
As a final perspective for our results, we envision two important developments: as for the case of the honeycomb lattice, an opportune modulation of the lattice periodicity can result in a homogeneous shift of the Fermi energy [18]. This property can use used to create an effective $pn$-junction where the super-Klein tunneling [43] could be observed. Secondly, to enhance the possibility of observing the effects of electron correlation in the flat band, we plan to investigate in future studies the implementation of the dice lattice in semiconducting electron simulator platforms based on InAs and InSb decorated with In or Cs adatoms [55, 56]. Finally, regarding the external magnetic field, an interesting outlook would be to study the strain-induced pseudomagnetic field in the case of next-nearest-neighbor term [57] since intensities of the order of 60 hundred Tesla have been achieved for the case the honeycomb lattice [18].
Figure 8: Electronic energy spectrum for the configuration with 24 CO molecules calculated with both cylindrical (empty square) and Gaussian (filled red-circles) potential.
@@ Acknowledgements.
Appendix A Model and Parameters
Here, we discuss an additional shape for the muffin-thin potential and the dependence of the parameters for the case of the cylindrical one.
It turns out that the muffin-tin model with cylindrical potentials approximates very well the experimental results by choosing the parameters a potential barrier $V_{0}=0.9\,\text{eV}$ and a diameter $D=0.6\,\text{nm}$ [18, 25, 22]. An alternative description of the CO molecules is a two-dimensional isotropic Gaussian potential:
$V_{\text{Gauss}}(\bm{r}-\bm{r}_{0})=V_{0}^{(G)}\,\text{e}^{-\left(\frac{\bm{r}-\bm{r}_{0}}{\sqrt{2}\,\alpha}\right)^{2}},$ (22)
with $V_{0}^{(G)}=1.8\,\text{eV}$ and $\alpha=0.15\,\text{nm}$, such that
[…]
where $n=1,2$ and $V$ is cylindrical potential. In particular, we can assume that the Gaussian vanishes fast enough within the unit cell and we require that the first and second momenta of the two potentials coincide.
Figure 9: Electronic energy spectrum for the configuration with 24 CO molecules calculated with both muffin-tin potential strength $V_{0}=0.9\,\text{eV}$ (empty square) and $V_{0}=1\,\text{eV}$ (filled red circles). The diameter potential is $D=0.6\,\text{nm}$ in both cases.
The Fourier components $V_{\bm{G}}$ are in the Gaussian case
[…]
$=(2\pi\alpha^{2}V_{0})\text{e}^{-\lvert\bm{G}\rvert^{2}\alpha^{2}/2}\,$ (24a)
and we found that the kind of potential does not affect significantly the nearly free electron model (in Figure 8 we report the case with 24 CO molecules).
Finally, by increasing the muffin-tin cylindrical potential strength $V_{0}$, we may expect higher values of the on-site energies and higher ratio $t/t^{\prime}$ between the nearest and next-nearest-neighbor hopping terms, which is a measure of the central band flatness (note that $V_{0}$ is tunable only by changing the CO molecule with one with different repulsive potential [24]). This is shown in Figure 9 for the 24 CO molecules case, where the potential has increased by about 10%. This results in a lower gap $\Delta$ and in an increase of the $t/t^{\prime}$ ratio by about 6%.